jumping height on Δερκσεν name of the planet in Latin letters take-off points velocity on earth gravitational acceleration on the planets surface lifting, upward acceleration (Archimedes) /3 /4 /3 /4 6 /20r-6.67 1011 m3/kg s2 #1 Strange planet The 4th intergalactic Olympic Games take place on the planet Δερκσεν in the year 2115, The gearth-9.81 m/s2 great grandchild of Dirk Nowitzki is famous for high jump: 2.34 m on the earth! On another planet a different height is possible as the planets mass is 3-1024 kg and the radius is 5400 km The atmosphere is enriched with SF6 such that the density reaches 5 kg/m3. All sportsmen are wearing a special comfortable suit filled with air at a density of 1.4 kg/m3 and a volume of 2.3 m3. What is the weight if the sportsmens mass is 85 kg?
Answer #1

The weight of an object is defined by force at which the object is pulled towards the center.

So, to find the weight of sportsmen whose mass is 85 kg. we need to find the acceleration due to gravity of the given planet.

Now,

Given planet's mass=M = 3*1024 kg

radius =r = 5400km= 5400 *103 m= 5.4 * 106m

Gamma=6.67 * 10-11 m3/kgs2

acceleration due to gravity is given by

「M gー

6.6710024 (5.4 106)2 11 J ? ms

g = 6.86ms-2

So, weight of the sportsmen = 85 * 6.86 N = 583N, which is the required solution.

This is when atmosphere weight is not considered.

Since the atmosphere has also some weight the overall force it exerts downwards will also change.

Now, mass of atmosphere displaced by the human = density of atmosphere * volume = 5kg/m3 * 2.3 m3 =11.5 kg

mass of the air that is used by human to counteract some of the atmospheric forces = density of air * volume = 1.4 kg/m3 * 2.3 m3 = 3.22 kg

So, overall mass of the human changes to = mass of human + mass of atmosphere - mass of air = 85 +11.5-3.22 kg =93.28kg

So, overall weight = 93.28 * 6.86 N= 640 N.

If any doubt feel free to comment